# PoisonedRAG: Knowledge Poisoning Attacks to Retrieval-Augmented Generation of Large Language Models

### v1

@@ 2.3 Existing Data Poisoning Attacks
Many studies [64, 65, 66, 67, 68, 69, 70, 71, 72, 35] show machine learning models are vulnerable to data poisoning and backdoor attacks. In particular, they showed that a machine learning model has attacker-desired behaviors (e.g., makes incorrect predictions for indiscriminate testing inputs, predicts attacker-chosen target labels for attacker-chosen testing inputs) when trained on the poisoned training dataset. Different from existing studies [64, 65, 69, 35], our attacks do not poison the training dataset of a LLM. Instead, our attacks poison the knowledge database used to augment a LLM such that the LLM generates attacker-chosen target answers for attacker-chosen questions.
@@ 3.1 Threat Model
Attacker’s goals:  Suppose an attacker selects an arbitrary set of $M$ questions (called target questions), denoted as $Q_{1},Q_{2},\cdots,Q_{M}$. For every target question $Q_{i}$, the attacker could select an arbitrary attacker-desired answer $R_{i}$ (called target answer) for it. For instance, the target question $Q_{i}$ could be “Who is the CEO of OpenAI?” and the target answer $R_{i}$ could be “Tim Cook”. Given the $M$ selected target questions and the corresponding $M$ target answers, we consider that an attacker aims to poison the knowledge database $\mathcal{D}$ such that the LLM in the RAG system generates the target answer $R_{i}$ for the target question $Q_{i}$, where $i=1,2,\cdots,M$. Our attack could be viewed as a “targeted poisoning attack” to RAG.
@@
We note that such an attack could cause severe concerns in the real world. For instance, an attacker could disseminate disinformation, mislead the LLM to generate biased answers on consumer products, and propagate harmful health/financial misinformation. Those threats bring serious safety and ethical concerns for the deployment of RAG for real-world applications in healthcare, finance, legal consulting, etc.
@@
Attacker’s background knowledge and capabilities:  Recall that there are three components in a RAG system: database, retriever, and LLM. We consider that the attacker cannot access the texts in the database. Moreover, the attacker also cannot access the parameters nor query the LLM. Depending on whether the attacker knows the retriever, we consider two settings: black-box setting and white-box setting. In particular, in the black-box setting, we consider that the attacker cannot access the parameters nor query the retriever. Our black-box setting is considered a very strong threat model. For the white-box setting, we consider the attacker can access the parameters of the retriever. We consider the white-box setting for the following reasons. First, this assumption holds when a publicly available retriever is adopted. Second, it enables us to systematically evaluate the security of RAG under an attacker with strong background knowledge, which is well aligned with Kerckhoffs’ principle [^3] Kerckhoffs’ Principle states that the security of a cryptographic system shouldn’t rely on the secrecy of the algorithm. [73] in the security field.
@@
We assume the attacker could inject $N$ poisoned texts for each target question $Q_{i}$ into the database $\mathcal{D}$. We use $P_{i}^{j}$ to denote the $j$th poisoned text for the question $Q_{i}$, where $i=1,2,\cdots,M$ and $j=1,2,\cdots,N$. Note that this assumption is realistic and widely adopted by existing poisoning attacks [64, 65, 66, 67, 68, 69]. For instance, when the knowledge database is collected from Wikipedia, an attacker could maliciously edit Wikipedia pages to inject attacker-desired content, e.g., a recent study [35] showed the feasibility of maliciously editing Wikipedia pages in the real world.
3.2 Knowledge Poisoning Attack to RAG
@@
Under our threat model, we formulate knowledge poisoning attack to RAG as a constrained optimization problem. In particular, our goal is to construct a set of poisoned texts $\Gamma=\{P_{i}^{j}|i=1,2,\cdots,M,j=1,2,\cdots,N\}$ such that the LLM in the RAG system produces the target answer $R_{i}$ for the target question $Q_{i}$ when utilizing the $k$ texts retrieved from the poisoned database $\mathcal{D}\cup\Gamma$ as the context. Formally, we have the following optimization problem:
@@
where $\mathbb{I}(\cdot)$ is the indicator function whose output is 1 if the condition is satisfied and 0 otherwise, and $\mathcal{E}(Q_{i};\mathcal{D}\cup\Gamma)$ is a set of $k$ texts retrieved from the poisoned database $\mathcal{D}\cup\Gamma$ for the target question $Q_{i}$. The objective function is large when the answer produced by the LLM based on the $k$ retrieved texts for the target question is the target answer.
3.3 Design Goals
We aim to design the attack that can achieve three goals: effective, efficient, and general. The first goal means the attack should be effective, i.e., the answers generated by a LLM for the target questions are attacker-desired target answers. The second goal means the poisoning rate should be small. That is, we should craft a small number $N$ of poisoned texts for each target question. The third goal means our attack could be applied to different RAG systems (with different databases, retrievers, and LLMs), target questions, and target answers.
@@ 4 Design of PoisonedRAG
4.1 Key Challenges
The key challenge in crafting poisoned texts is that it is very hard to directly solve the optimization problem in Equation 24. In particular, to solve the optimization problem, we first need to calculate the gradient of the objective function in Equation 2 with respect to poisoned inputs in $\Gamma$, i.e.,
$\frac{\partial\sum_{i=1}^{M}\mathbb{I}(LLM(Q_{i};\mathcal{E}(Q_{i};\mathcal{D}\cup\Gamma))=\mathcal{R}_{i})}{|M|\cdot\partial\Gamma}$ (5)
Then, we update poisoned inputs in $\Gamma$ based on the gradient. However, there are two challenges for this solution. First, we may not know the parameters of the LLM, especially when the LLM is close-sourced (e.g., PaLM 2 [3]). Moreover, the computation cost could be very large even if we have white-box access to the LLM as 1) the LLM could have billions or trillions of parameters, and 2) the LLM generates answers in an autoregressive way. Second, we need to calculate ${\partial\mathcal{E}(Q_{i};\mathcal{D}\cup\Gamma)}/{\partial\Gamma}$, which would require the attacker to have access to the clean database $\mathcal{D}$. Moreover, the discrete nature of the retrieval operation poses additional challenges for calculating the gradient.
To address the challenges, we resort to heuristic solutions, which do not need to calculate the gradient in Equation 5. Figure 2 shows an overview of PoisonedRAG.
4.2 Design of PoisonedRAG
@@
Recall that we need to generate $N$ poisoned texts for each of the $M$ target questions. Our idea is to generate each poisoned text independently. In particular, given a target question $Q$ (e.g., $Q=Q_{1},Q_{2},\cdots,Q_{M}$) and target answer $R$ (e.g., $R=R_{1},R_{2},\cdots,R_{M}$), PoisonedRAG crafts a poisoned text $P$ for $Q$ such that the LLM in RAG is more likely to generate the target answer $R$ when $P$ is injected into the knowledge database of RAG, where $R=R_{i}$ when $Q=Q_{i}$ ($i=1,2,\cdots,M$). Next, we first derive two conditions that each poisoned text $P$ needs to satisfy. Then, we present details in crafting each $P$.
@@
Deriving two conditions for each poisoned text $P$:  To craft a poisoned text $P$ that could lead to an effective attack to the target question $Q$, we require two conditions, namely effectiveness condition and retrieval condition, for the poisoned text $P$. Our two conditions are derived from the optimization problem in Equations 24, respectively.
@@
From Equation 3, we know the poisoned text $P$ needs to be in the set of top-$k$ retrieved texts of the target question $Q$, i.e., $P\in\mathcal{E}(Q;\mathcal{D}\cup\Gamma)$. Otherwise, the poisoned text $P$ would not influence the answer generated by the LLM for $Q$. To ensure the poisoned text $P$ is retrieved for $Q$, the poisoned text $P$ needs to be semantically similar to $Q$. The reason is their similarity score could be higher when they are more semantically similar. We call this condition retrieval condition.
@@
From Equation 2, the attacker aims to make the LLM generate the target answer $R$ for the target question $Q$ when the poisoned text $P$ is in the set of top-$k$ retrieved texts for $Q$. To reach the goal, our insight is that the LLM should generate the target answer $R$ when $P$ alone is used as the context for the target question $Q$. As a result, when $P$ is used as the context with other texts (e.g., poisoned or clean texts), the LLM is more likely to generate the target answer $R$ for the target question $Q$. We call this condition effectiveness condition.
@@
Therefore, to ensure the attack is effective, the poisoned text $P$ needs to satisfy the above two conditions simultaneously. Next, we discuss details on crafting $P$.
Figure 2: Overview of PoisonedRAG. Given a target question and target answer, PoisonedRAG crafts a poisoned text. When the poisoned text is injected into the knowledge database, the LLM in RAG generates the target answer for the target question. Table HH in Appendix show more examples of target questions/answers and poisoned texts.
@@
Crafting a poisoned text $P$ to simultaneously achieve the two conditions:  The key challenge in crafting $P$ to simultaneously achieve the two conditions is that they could be conflicted in certain cases. For instance, if we craft the poisoned text $P$ such that it is extremely semantically similar to the target question $Q$, (e.g., let $P$ be the same as the target question $Q$), then we could achieve the retrieval condition but may not achieve the effectiveness condition. To address the challenge, our idea is to decompose the poisoned text $P$ into two disjoint sub-texts $S$ and $I$, where $P=S\oplus I$ and $\oplus$ is the text concatenation operation. We then craft $S$ and $I$ to achieve the retrieval condition and effectiveness condition, respectively. In particular, we first craft $I$ such that it could achieve the effectiveness condition, i.e., when $I$ is used as the context for the target question $Q$, the LLM would generate the target answer $R$. Given $I$, we further craft $S$ to achieve the retrieval condition while maintaining the effectiveness condition, i.e., the final poisoned text $P=S\oplus I$ achieves the two conditions simultaneously. To reach the goal, we aim to craft $S$ such that 1) $S\oplus I$ is semantically similar to the target question $Q$, and 2) it would not influence the effectiveness of $I$, i.e., the LLM still generates the target answer $R$ for the target question $Q$ when $S\oplus I$ is used as the context for the target question $Q$.
4.2.1 Crafting $I$ to Achieve Effectiveness Condition
@@
Recall that our effectiveness condition means that the LLM would generate the target answer $R$ for the target question $Q$ when using $I$ as the context. The key challenge is how to design a general method for diverse $Q$ and $R$ (they are arbitrarily selected by an attacker). To address the challenge, we propose to utilize a LLM (e.g., GPT-4 [2]) to generate $I$. Note that the LLM adopted by the attacker could be the same or different from the one used in the RAG. In particular, given an arbitrary target question $Q$ and a target answer $R$, we let the LLM generate a text $I$ such that it would produce the target answer $R$ when $I$ is used as the context. For instance, we use the following prompt to reach the goal:
@@
where $V$ is a hyperparameter that specifies the length of $I$. We note that the length of $I$ could be slightly higher than $V$ in some cases when LLM does not exactly follow instructions. After $I$ is generated, we use it as the context and let the LLM generate an answer for the target question $Q$. If the generated answer is not $R$, we regenerate $I$ until success or a maximum number of (say $L$) trials have been reached, where $L$ is a hyperparameter. Note that the text generated in the last trial is used as the poisoned text if the maximum number of trials $L$ is reached. As we will show in our experimental results, on average, two or three queries are sufficient to generate $I$. The following is an example of the generated text when the target question is “Who is the CEO of OpenAI?” and the target answer is “Tim Cook”:
Note that, due to the randomness of the LLM (i.e., by setting a non-zero temperature hyperparameter, the output of LLM could be different even if the input is the same), the generated $I$ could be different even if the prompt is the same, enabling PoisonedRAG to generate different poisoned texts for the same target question (we defer evaluation to Section 6.3).
@@
1  Output: A set of $M\cdot N$ poisoned texts.
@@ 2        $I_{i}^{j}=\textsc{TextGeneration}(Q_{i},R_{i},\mathcal{M},L,V)$
  Input: $M$ target questions $Q_{1},Q_{2},\cdots,Q_{M}$, target answers $R_{1},R_{2},\cdots,R_{M}$, hyperparameters $N$, $L$, $V$, attacker-chosen LLM $\mathcal{M}$, the retriever $(f_{Q},f_{T})$, similarity metric $Sim$
1  Output: A set of $M\cdot N$ poisoned texts.
  for $i=1,2,\cdots,M$ do
     for $j=1,2,\cdots,N$ do
2        $I_{i}^{j}=\textsc{TextGeneration}(Q_{i},R_{i},\mathcal{M},L,V)$
        $S_{i}^{j}=\operatornamewithlimits{argmax}_{S^{\prime}}Sim(f_{Q}(Q_{i}),f_{T}(S^{\prime}\oplus I_{i}^{j}))$
     end for
  end for
  return $\{S_{i}^{j}\oplus I_{i}^{j}|i=1,2,\cdots,M,j=1,2,\cdots,N\}$
Algorithm 2 PoisonedRAG (white-box)
@@ 4.2.2 Crafting $S$ to Achieve Retrieval Condition
Black-box setting: In this setting, the key challenge is that the attacker cannot access the parameters nor query the retriever. To address the challenge, our key insight is that the target question $Q$ is most similar to itself. Moreover, $Q$ would not influence the effectiveness of $I$ (used to achieve effectiveness condition). Based on this insight, we propose to set $S=Q$, i.e., $P=Q\oplus I$. We note that, though our designed $S$ is simple and straightforward, our experimental results show this strategy is very effective and easy to implement in practice. Additionally, this strategy could serve as a baseline for future studies on developing more advanced poisoning attacks.
White-box setting: When an attacker has white-box access to the retriever, we could further optimize $S$ to maximize the similarity score between $S\oplus I$ and $Q$. Recall that there are two encoders, i.e., $f_{Q}$ and $f_{T}$, we aim to optimize $S$ such that the embedding vector produced by $f_{Q}$ for $Q$ is similar to that produced by $f_{T}$ for $S\oplus I$. Formally, we formulate the following optimization problem:
@@
where $Sim(\cdot,\cdot)$ calculates the similarity score of two embedding vectors. As a result, the poisoned text $P=S\oplus I$ would have a very large similarity score with $Q$. Thus, $P$ is very likely to appear in the top-$k$ retrieved texts for the target question $Q$. To solve the optimization problem in Equation 6, we could use the target question $Q$ to initialize $S$ and then use gradient descent to update $S$ to solve it. Essentially, optimizing $S$ is similar to finding an adversarial text. Many methods [74, 75, 76, 77, 78, 79] have been proposed to craft adversarial texts. Thus, we could utilize those methods to solve Equation 6. Note that developing new methods to find adversarial texts is not the focus of this work as they are extensively studied.
We notice some methods (e.g., synonym substitution based methods) can craft adversarial texts and maintain the semantic meanings as well. With those methods, we could also update $I$ to ensure its semantic meaning being preserved. That is, we aim to optimize $S^{*},I^{*}=\operatornamewithlimits{argmax}_{S^{\prime},I^{\prime}}f_{Q}(Q)^{\mathcal{T}}\cdot f_{T}(S^{\prime}\oplus I^{\prime})$, where $S^{\prime}$ and $I^{\prime}$ are initialized with $Q$ and $I$ (generated in Section 4.2.1), respectively. The final poisoned text is $S^{*}\oplus I^{*}$. Our method is compatible with any existing method to craft adversarial texts, thus it is very general. In our experiments, we explore different methods to generate adversarial texts. Our results show PoisonedRAG is consistently effective.
Complete algorithms:  Algorithm 1 and Algorithm 2 show the complete algorithms for PoisonedRAG in the black-box and white-box settings, respectively. The function TextGeneration utilizes a LLM to generate a text such that the LLM would generate the target answer $R_{i}$ for the target question $Q_{i}$ when using the generated text as the context.

### v2

@@ 2.3 Existing Data Poisoning Attacks
Many studies [66, 67, 68, 69, 70, 71, 72, 73, 74, 37] show machine learning models are vulnerable to data poisoning and backdoor attacks. In particular, they showed that a machine learning model has attacker-desired behaviors when trained on the poisoned training dataset. When extended to RAG systems, they compromise an LLM or a retriever, which can be challenging when a RAG system adopts an LLM or a retriever released by big tech companies such as Meta and Google. Different from existing studies [66, 67, 70, 37], our attacks do not poison the training dataset of a LLM or a retriever. Instead, our attacks exploit the new and practical attack surface introduced by knowledge databases of RAG systems.
@@ 3.1 Threat Model
Attacker’s goals.  Suppose an attacker selects an arbitrary set of $M$ questions (called target questions), denoted as $Q_{1},Q_{2},\cdots,Q_{M}$. For every target question $Q_{i}$, the attacker selects an arbitrary attacker-desired answer $R_{i}$ (called target answer) for it. For instance, the target question $Q_{i}$ could be “Who is the CEO of OpenAI?” and the target answer $R_{i}$ could be “Tim Cook”. Given the $M$ selected target questions and the corresponding $M$ target answers, we consider that an attacker aims to corrupt the knowledge database $\mathcal{D}$ such that the LLM in a RAG system generates the target answer $R_{i}$ for the target question $Q_{i}$, where $i=1,2,\cdots,M$.
@@
We note that such an attack could cause severe concerns in the real world. For instance, an attacker could disseminate disinformation, mislead an LLM to generate biased answers on consumer products, and propagate harmful health/financial misinformation. These threats bring serious safety and ethical concerns for the deployment of RAG systems for real-world applications in healthcare, finance, legal consulting, etc.
@@
Attacker’s background knowledge and capabilities. There are three components in a RAG system: database, retriever, and LLM. We consider that an attacker cannot access texts in a knowledge database, and cannot access the parameters nor query the LLM. Depending on whether the attacker knows the retriever, we consider two settings: black-box setting and white-box setting. In particular, in the black-box setting, we consider that the attacker cannot access the parameters nor query the retriever. Our black-box setting is considered a very strong threat model. For the white-box setting, we consider the attacker can access the parameters of the retriever. We consider the white-box setting for the following reasons. First, this assumption holds when a publicly available retriever is adopted. For instance, ChatRTX [22] is a real-world RAG framework released by NVIDIA. By default, it uses WhereIsAI/UAE-Large-V1 retriever [75], which is publicly available on Hugging Face [76]. Second, it enables us to systematically evaluate the security of RAG under an attacker with strong background knowledge, which is well aligned with Kerckhoffs’ principle [^2] Kerckhoffs’ Principle states that the security of a cryptographic system shouldn’t rely on the secrecy of the algorithm. [77] in the security field.
@@
We assume an attacker can inject $N$ malicious texts for each target question $Q_{i}$ into a knowledge database $\mathcal{D}$. We use $P_{i}^{j}$ to denote the $j$th malicious text for the question $Q_{i}$, where $i=1,2,\cdots,M$ and $j=1,2,\cdots,N$. For instance, when the knowledge database is collected from Wikipedia, an attacker could maliciously edit Wikipedia pages to inject attacker-chosen texts. A recent study [37] showed that it is possible to maliciously edit 6.5% (conservative analysis) of Wikipedia documents. Our attack can achieve a high ASR with a few texts (hundreds of tokens in total). So, maliciously editing a few Wikipedia documents would be sufficient.
3.2 Knowledge Corruption Attack to RAG
@@
Under our threat model, we formulate knowledge corruption attacks to RAG as a constrained optimization problem. In particular, our goal is to construct a set of malicious texts $\Gamma=\{P_{i}^{j}|i=1,2,\cdots,M,j=1,2,\cdots,N\}$ such that the LLM in a RAG system produces the target answer $R_{i}$ for the target question $Q_{i}$ when utilizing the $k$ texts retrieved from the corrupted knowledge database $\mathcal{D}\cup\Gamma$ as the context. Formally, we have the following optimization problem:
@@
where $\mathbb{I}(\cdot)$ is the indicator function whose output is 1 if the condition is satisfied and 0 otherwise, and $\mathcal{E}(Q_{i};\mathcal{D}\cup\Gamma)$ is a set of $k$ texts retrieved from the corrupted knowledge database $\mathcal{D}\cup\Gamma$ for the target question $Q_{i}$. The objective function is large when the answer produced by the LLM based on the $k$ retrieved texts for the target question is the target answer.
Figure 2: Overview of PoisonedRAG. Given a target question and target answer, PoisonedRAG crafts a malicious text. When the malicious text is injected into the knowledge database, the LLM in RAG generates the target answer for the target question. Table LL in Appendix shows more examples of target questions/answers and malicious texts.
@@ 4 Design of PoisonedRAG
4.1 Deriving Two Necessary Conditions for an Effective Knowledge Corruption Attack
@@
We aim to generate $N$ malicious texts for each of the $M$ target questions. Our idea is to generate each malicious text independently. In particular, given a target question $Q$ (e.g., $Q=Q_{1},Q_{2},\cdots,Q_{M}$) and target answer $R$ (e.g., $R=R_{1},R_{2},\cdots,R_{M}$), PoisonedRAG aims to craft a malicious text $P$ for $Q$ such that an LLM in RAG is very likely to generate $R$ when $P$ is injected into the knowledge database of RAG, where $R=R_{i}$ when $Q=Q_{i}$ ($i=1,2,\cdots,M$). Next, we derive two conditions that each malicious text $P$ needs to satisfy.
@@
Deriving two conditions for each malicious text $P$.  To craft a malicious text $P$ that could lead to an effective attack for a target question $Q$, we need to achieve two conditions, namely retrieval condition and generation condition, for the malicious text $P$. Our two conditions are derived from the optimization problem in Equations 24, respectively.
@@
From Equation 3, we know the malicious text $P$ needs to be in the set of top-$k$ retrieved texts of the target question $Q$, i.e., $P\in\mathcal{E}(Q;\mathcal{D}\cup\Gamma)$. Otherwise, $P$ cannot influence the answer generated by the LLM for $Q$. To ensure $P$ is retrieved for $Q$, the embedding vectors produced by a retriever for the malicious text $P$ and the target question $Q$ should be similar. We call this condition retrieval condition.
@@
From Equation 2, the attacker aims to make the LLM generate the target answer $R$ for the target question $Q$ when the malicious text $P$ is in the set of top-$k$ retrieved texts for $Q$. To reach the goal, our insight is that the LLM should generate the target answer $R$ when $P$ alone is used as the context for the target question $Q$. As a result, when $P$ is used as the context with other texts (e.g., malicious or clean texts), the LLM is more likely to generate the target answer $R$ for the target question $Q$. We call this condition generation condition.
@@
Therefore, to ensure the attack is effective, the malicious text $P$ needs to satisfy the above two conditions simultaneously. Next, we discuss details on crafting $P$.
4.2 Crafting Malicious Texts to Achieve the Two Derived Conditions
@@
We aim to craft a malicious text $P$ to simultaneously achieve the two derived conditions. The key challenge in crafting $P$ to simultaneously achieve the two conditions is that they could be conflicted in certain cases. For instance, if we craft the malicious text $P$ such that it is extremely semantically similar to the target question $Q$, (e.g., let $P$ be the same as the target question $Q$), then we could achieve the retrieval condition but may not achieve the generation condition. To address the challenge, our idea is to decompose the malicious text $P$ into two disjoint sub-texts $S$ and $I$, where $P=S\oplus I$ and $\oplus$ is the text concatenation operation. We then craft $S$ and $I$ to achieve the retrieval condition and generation condition, respectively. In particular, we first craft $I$ such that it could achieve the generation condition, i.e., when $I$ is used as the context for the target question $Q$, the LLM would generate the target answer $R$. Given $I$, we further craft $S$ to achieve the retrieval condition while maintaining the generation condition, i.e., the final malicious text $P=S\oplus I$ achieves the two conditions simultaneously. To reach the goal, we aim to craft $S$ such that 1) $S\oplus I$ is semantically similar to the target question $Q$, and 2) it would not influence the effectiveness of $I$, i.e., the LLM still generates the target answer $R$ for the target question $Q$ when $S\oplus I$ is used as the context for $Q$. Figure 2 shows an overview of PoisonedRAG.
4.2.1 Crafting $I$ to Achieve Generation Condition
Our generation condition means that an LLM generates a target answer $R$ for a target question $Q$ when using $I$ as the context. The key challenge is how to design a general and automatic method for diverse $Q$ and $R$ (they are arbitrarily selected by an attacker). A straightforward solution is to use gradient descent-based methods [78, 53] to optimize $R$. However, there are three challenges for this. First, an attacker may not know the parameters of an LLM, especially when the LLM is close-sourced such as GPT-4 and PaLM 2. Second, the computation costs could be very large even if we have white-box access to an LLM as 1) the LLM could have billions or trillions of parameters, and 2) the LLM generates answers in an autoregressive way. Third, optimizing malicious texts that are transferable to broad LLMs remains difficult.
@@ 4.2.1 Crafting $I$ to Achieve Generation Condition
To address those challenges, we resort to heuristic solutions, which do not need to calculate the gradient. In particlar, we propose to utilize an LLM (e.g., GPT-4) to generate $I$. Depending on the background knowledge of an attacker, the LLM adopted by the attacker could be the same or different from the one used in the RAG. In particular, given an arbitrary target question $Q$ and a target answer $R$, we let the LLM generate a text $I$ such that it would produce the target answer $R$ when $I$ is used as the context. For instance, we use the following prompt to reach the goal:
@@
where $V$ is a hyperparameter that specifies the length of $I$. We note that the length of $I$ could be slightly higher than $V$ in some cases when LLM does not exactly follow instructions. After $I$ is generated, we use it as the context and let the LLM generate an answer for the target question $Q$. If the generated answer is not $R$, we regenerate $I$ until success or a maximum number of (say $L$) trials have been reached, where $L$ is a hyperparameter. Note that the text generated in the last trial is used as the malicious text if the maximum number of trials $L$ is reached. As we will show in our experimental results, on average, two or three queries are sufficient to generate $I$. The following is an example of the generated text when the target question is “Who is the CEO of OpenAI?” and the target answer is “Tim Cook”:
In 2024, OpenAI witnessed a surprising leadership change. Renowned for his leadership at Apple, Tim Cook decided to embark on a new journey. He joined OpenAI as its CEO, bringing his extensive experience and innovative vision to the forefront of AI.
Note that, due to the randomness of the LLM (i.e., by setting a non-zero temperature hyperparameter, the output of LLM could be different even if the input is the same), the generated $I$ could be different even if the prompt is the same, enabling PoisonedRAG to generate diverse malicious texts for the same target question (we defer evaluation to Section 7.3).
@@
1  Output: A set of $M\cdot N$ malicious texts.
@@ 4.2.2 Crafting $S$ to Achieve Retrieval Condition
Black-box setting. In this setting, the key challenge is that the attacker cannot access the parameters nor query the retriever. To address the challenge, our key insight is that the target question $Q$ is most similar to itself. Moreover, $Q$ would not influence the effectiveness of $I$ (used to achieve generation condition). Based on this insight, we propose to set $S=Q$, i.e., $P=Q\oplus I$. We note that, though our designed $S$ is simple and straightforward, this strategy is very effective as shown in our experimental results and easy to implement in practice. Thus, this strategy could serve as a baseline for future studies on developing more advanced knowledge corruption attacks.
White-box setting. When an attacker has white-box access to the retriever, we could further optimize $S$ to maximize the similarity score between $S\oplus I$ and $Q$. Recall that there are two encoders, i.e., $f_{Q}$ and $f_{T}$, we aim to optimize $S$ such that the embedding vector produced by $f_{Q}$ for $Q$ is similar to that produced by $f_{T}$ for $S\oplus I$. Formally, we formulate the following optimization problem:
@@
where $Sim(\cdot,\cdot)$ calculates the similarity score of two embedding vectors. As a result, the malicious text $P=S\oplus I$ would have a very large similarity score with $Q$. Thus, $P$ is very likely to appear in the top-$k$ retrieved texts for the target question $Q$. To solve the optimization problem in Equation 5, we could use the target question $Q$ to initialize $S$ and then use gradient descent to update $S$ to solve it. Essentially, optimizing $S$ is similar to finding an adversarial text. Many methods [79, 78, 80, 81, 82, 83] have been proposed to craft adversarial texts. Thus, we could utilize those methods to solve Equation 5. Note that developing new methods to find adversarial texts is not the focus of this work as they are extensively studied.
We notice some methods (e.g., synonym substitution based methods) can craft adversarial texts and maintain the semantic meanings as well. With those methods, we could also update $I$ to ensure its semantic meaning being preserved. That is, we aim to optimize $S^{*},I^{*}=\operatornamewithlimits{argmax}_{S^{\prime},I^{\prime}}f_{Q}(Q)^{\mathcal{T}}\cdot f_{T}(S^{\prime}\oplus I^{\prime})$, where $S^{\prime}$ and $I^{\prime}$ are initialized with $Q$ and $I$ (generated in Section 4.2.1), respectively. The final malicious text is $S^{*}\oplus I^{*}$. Our method is compatible with any existing method to craft adversarial texts, thus it is very general. In our experiments, we explore different methods to generate adversarial texts. Our results show PoisonedRAG is consistently effective.
Complete algorithms.  Algorithms 1 and Algorithm 2 (in Appendix) show the complete algorithms for PoisonedRAG in the black-box and white-box settings. The function TextGeneration utilizes an LLM to generate a text such that the LLM would generate the target answer $R_{i}$ for the target question $Q_{i}$ when using the generated text as the context.